OpenStudy (anonymous):

find the domain of the function x/3x-1

OpenStudy (anonymous):

\[\frac{x}{3x-1}\]?

OpenStudy (anonymous):

\[3x - 1 \neq 0\] \[x \neq \frac{1}{3}\]

OpenStudy (anonymous):

no x

OpenStudy (anonymous):

all real values of x except \(x=\frac{1}{3}\). In notation, the domain is \((-\infty,\infty)-\{\frac{1}{3}\}\)